What does partitioning mean? Explain and illustrate your answer.


1
Expert's answer
2021-05-31T18:30:29-0400

Partitioning is a tactic of solving mathematical problems involving large numbers by splitting them into small units to enable easy computation.

For example,

79+34=113


The minors are taught to split it into

70+9+30+4=113


Or 70+30+9+4=113